OpenStudy (anonymous):

what is the mass of hydrogen is formed when 3 mol of aluminum react with excess hydrochloric acid 2Al + 6HCl -->2AlCl3 + 3H2

OpenStudy (aaronq):

use the coefficients and the moles (n) of each species in a ratio, like this: For a general equation: aA +bB -> cC where a,b,c (lowercase) are the coefficients and A,B,C are the species \[\frac{ n _{A} }{ a }=\frac{ n _{B} }{ b }=\frac{ n _{C}}{ c }\]

OpenStudy (aaronq):

then convert moles of H2 to mass using moles=mass/molar mass

OpenStudy (anonymous):

Hey Thanks for the post! I am still not clear on what I should be doing. This is what I have /understood from your post: I calculated the molar mass of each species and divide it by its coefficient.... 26.9815/2=13.49, 36.4609/6=6.07681, 133.3405/2=66.6702 Im not sure what I should do next

OpenStudy (aaronq):

it's moles not molar mass :P you have to know that in chemical reactions you always use moles (except when you use gases, you can take shortcuts). you've already been given moles so theres not need to convert anything. so, you have: 2Al + 6HCl -->2AlCl3 + 3H2 you're given 3 moles of Al to react and wanna find how much H2 was produced. the equation relating them is set up like this: \[\frac{ n _{Al} }{ 2 }=\frac{ n _{H _{2}} }{ 3 }\] now you plug your values into it and solve algebraically.

OpenStudy (aaronq):

n = moles, btw.

OpenStudy (anonymous):

Thanks for the help and tips! You really helped me out! :)